OpenStudy (anonymous):

it takes 226 feet of fence to surrond the rectangular park in Jackson Heights.Randy Watson recalls that when the park was built the contractor made sure that the length was 11 less than three times the width. he did this because his son was 3 years old and was born on the 11th of the month. what are the dimensions of the park in Jackson Heights?

OpenStudy (anonymous):

so first make an equation, you know that the perimeter of the rectangle is 226ft and the relationship of the length ( L ) to the width ( w ) is w = 3L - 11 you also know that 2L + 2w = 226 with these you can solve for each variable plug the first equation into the second in place of w leaving you with: 2L + 2 (3L - 11) = 226 divide by 2 L + 3L -11 = 113 add 11 to both sides and simplify L + 3L 4L = 124 divide by 4 L = 31 now plug this into the other equation for w w = 3 (31) - 11 w = 91 - 11 w = 80 there you go, you now have both the length (L) and width (w)! :D

OpenStudy (anonymous):

Thank Very Much:D Can You Please Help Me With Another One?

OpenStudy (anonymous):

sure